OpenStudy (anonymous):

What is the solution of the equation? \[(x -2)^{3/4} = 8\]

OpenStudy (anonymous):

((x-2)^(3/4))^(4/3)=8^(4/3) x-2=16 ......

OpenStudy (anonymous):

x=18